# Aligning Large Language Models to Low-Resource Languages through LLM-Based Selective Translation: A Systematic Study

### v1

@@ Abstract
Multilingual large language models (LLMs) often demonstrate a performance gap between English and non-English languages, particularly in low-resource settings. Aligning these models to low-resource languages is essential yet challenging due to limited high-quality data. While English alignment datasets are readily available, curating equivalent data in other languages is expensive and time-consuming. A common workaround is to translate existing English alignment data; however, standard translation techniques often fail to preserve critical elements such as code, mathematical expressions, and structured formats like JSON. In this work, we investigate LLM-based selective translation, a technique that selectively translates only the translatable parts of a text while preserving non-translatable content and sentence structure. We conduct a systematic study to explore key questions around this approach, including its effectiveness compared to vanilla translation, the importance of filtering noisy outputs, and the benefits of mixing translated samples with original English data during alignment. Our experiments focus on the low-resource Indic language Hindi and compare translations generated by Google Cloud Translation (GCP) and Llama-3.1-405B. The results highlight the promise of selective translation as a practical and effective method for improving multilingual alignment in LLMs.
@@ 1 Introduction
  • How does LLM-based selective translation of alignment data compare to conventional (vanilla) translation methods such as Google Cloud Translation (GCP) on the performance of the aligned model?
@@
We train the Nemotron-4-Mini-Hindi-4B-Base [^1] https://huggingface.co/nvidia/Nemotron-4-Mini-Hindi-4B-Base model on the LLM-translated and GCP-translated datasets and compare the performance on the downstream tasks like MTBench, IFEval, and GSM8K in Hindi. Our experiments focus on Hindi, a widely spoken yet low-resource Indic language. We compare translation quality and alignment effectiveness across outputs generated by GCP and Llama-3.1-405B, a powerful open-source LLM. Through this comprehensive analysis, we demonstrate that LLM-based selective translation offers a practical and robust solution for multilingual alignment, substantially improving the performance of LLMs in low-resource settings and moving toward more linguistically inclusive AI systems. Throughout this study, "LLM translation" specifically refers to translations produced by Llama-3.1-405B.
@@ 3.5 Experimental Design
  • GCP vs Llama-3.1-405B Translation: This experiment empirically compares the performance of models aligned using data translated by Google Cloud Translation (GCP) against those aligned with data generated via Llama-3.1-405B based selective translation. To achieve this, we perform SFT on the Nemotron-4-Mini-Hindi-4B-Base model. The SFT process utilizes a fixed set of 200k English data samples, combined with varying subsets of translated + filtered Hindi data. The Hindi filtered data subsets consist of 20K, 40K, 60K, 80K, and 100K samples, each randomly selected from a pool of 100K filtered examples. The unfiltered subset comprises 200K samples. Following SFT, the fine-tuned models are benchmarked for their performance across various Hindi test sets, including MTBench, IFEval, and GSM8K, to provide a comprehensive comparison. The Llama-3.1-405B was selected for this study as it represents the largest and highest-quality LLM available for Hindi translations, with human annotators consistently rating its translation quality superior to other competitor models like Llama-3.3-70B-Instruct and Nemotron-4-340B-Instruct. We note that while larger contexts can sometimes lead to "sentence-drop" issues in translation, this is not a concern here as we translate entire prompts and responses, ensuring coherence. Furthermore, larger LLMs like Llama-3.1-405B exhibit greater resilience to the sentence drop issue and a larger context length of 128k.
@@ 3.6 Evaluation Datasets
The evaluation of conversational abilities in large language models typically relies on extensive English datasets like IFEval, MTBench, and GSM8K. For Hindi, however, available options such as MILU Verma et al. (2024), and Global MMLU Singh et al. (2024) are more limited, primarily focusing on foundational model assessment rather than advanced conversational nuances. Direct translation of English datasets into Hindi often overlooks cultural nuances and linguistic structures, leading to grammatical errors and compounding inherent errors in the translation process. To address this, we adopt a multi-step approach incorporating human oversight to ensure accurate assessment of Hindi language capabilities. The following datasets were internally curated to benchmark the aligned models trained in this work.
@@
  • IFEval-Hi: The Hindi IFEval dataset contains 848 prompts to evaluate the instruction-following ability of LLMs in Hindi. Structured similarly to its English counterpart, it features "verifiable instructions" with heuristically validated responses. These prompts are natively curated by Hindi-proficient specialists to capture local linguistic nuances and Indian cultural context.
  • GSM8k-Hi: Hindi-GSM8K is the GCP-translated version of the English GSM8K test set. Its samples are meticulously reviewed and corrected by human annotators for quality improvement. These problems typically require 2 to 8 steps to solve, primarily involving elementary arithmetic calculations.
  • MTBench-Hi: The Hindi MTBench dataset consists of 200 multi-turn prompts designed to evaluate the conversational ability of Hindi LLMs. Eighty percent of its samples are natively created by Hindi specialists, with the remaining 20% translated from the English version, ensuring a balanced and comprehensive evaluation. For this work, we use a subset of 40 samples from MTBench-Hi, focusing on classes such as coding, STEM, math, reasoning, and multiturn interactions. The evaluation is conducted by GPT-4o, with responses rated on a scale of 1-10. It is noted that the scores obtained are on the lower side, as this subset represents areas where Hindi models typically do not excel.
@@ Appendix A Appendix
I have the following sentences:

### v2

@@ Abstract
Multilingual large language models (LLMs) often demonstrate a performance gap between English and non-English languages, particularly in low-resource settings. Aligning these models to low-resource languages is essential yet challenging due to limited high-quality data. While English alignment datasets are readily available, curating equivalent data in other languages is expensive and time-consuming. A common workaround is to translate existing English alignment data; however, standard translation techniques often fail to preserve critical elements such as code, mathematical expressions, and structured formats like JSON. In this work, we investigate LLM-based selective translation, a technique that selectively translates only the translatable parts of a text while preserving non-translatable content and sentence structure. We conduct a systematic study to explore key questions around this approach, including its effectiveness compared to vanilla translation, the importance of filtering noisy outputs, and the benefits of mixing translated samples with original English data during alignment. Our experiments focus on the low-resource Indic language Hindi and compare translations generated by Google Cloud Platform (GCP) and Llama-3.1-405B. The results highlight the promise of selective translation as a practical and effective method for improving multilingual alignment in LLMs.
@@ 1 Introduction
  • How does LLM-based selective translation of alignment data compare to conventional (vanilla) translation methods, such as Google Cloud Platform (GCP), on the performance of the aligned model?
@@
We train the Nemotron-4-Mini-Hindi-4B-Base [^1] https://huggingface.co/nvidia/Nemotron-4-Mini-Hindi-4B-Base Joshi et al. (2024) model on the LLM-translated and GCP-translated datasets and compare the performance on the downstream tasks like MTBench, IFEval, and GSM8K in Hindi. Our experiments focus on Hindi, a widely spoken yet low-resource Indic language. We compare translation quality and alignment effectiveness across outputs generated by GCP and Llama-3.1-405B, a powerful open-source LLM. Through this comprehensive analysis, we demonstrate that LLM-based selective translation offers a practical and robust solution for multilingual alignment, substantially improving the performance of LLMs in low-resource settings and moving toward more linguistically inclusive AI systems. Throughout this study, "LLM translation" specifically refers to translations produced by Llama-3.1-405B.
@@ 3.5 Experimental Design
  • GCP vs Llama-3.1-405B Translation: This experiment empirically compares the performance of models aligned using data translated by Google Cloud Platform (GCP) against those aligned with data generated via Llama-3.1-405B based selective translation. To achieve this, we perform SFT on the Nemotron-4-Mini-Hindi-4B-Base model. The SFT process utilizes a fixed set of 200k English data samples, combined with varying subsets of translated + filtered Hindi data. The Hindi filtered data subsets consist of 20K, 40K, 60K, 80K, and 100K samples, each randomly selected from a pool of 100K filtered examples. The unfiltered subset comprises 200K samples. Following SFT, the fine-tuned models are benchmarked for their performance across various Hindi test sets, including MTBench, IFEval, and GSM8K, to provide a comprehensive comparison. The Llama-3.1-405B was selected for this study as it represents the largest and highest-quality LLM available for Hindi translations, with human annotators consistently rating its translation quality superior to other competitor models like Llama-3.3-70B-Instruct and Nemotron-4-340B-Instruct. We note that while larger contexts can sometimes lead to "sentence-drop" issues in translation, this is not a concern here as we translate entire prompts and responses, ensuring coherence. Furthermore, larger LLMs like Llama-3.1-405B exhibit greater resilience to the sentence drop issue and a larger context length of 128k.
@@ 3.6 Evaluation Datasets
The evaluation of conversational abilities in large language models typically relies on extensive English datasets like IFEval, MTBench, and GSM8K. For Hindi, however, available options such as MILU Verma et al. (2024), and Global MMLU Singh et al. (2024) are more limited, primarily focusing on foundational model assessment rather than advanced conversational nuances. Direct translation of English datasets into Hindi often overlooks cultural nuances and linguistic structures, leading to grammatical errors and compounding inherent errors in the translation process. To address this, we adopt a multi-step approach incorporating human oversight to ensure accurate assessment of Hindi language capabilities. The following datasets introduced in Kamath et al. (2025) were used to benchmark the aligned models trained in this work.
@@
  • IFEval-Hi [^3] https://huggingface.co/datasets/nvidia/IFEval-Hi: The Hindi IFEval dataset contains 848 prompts to evaluate the instruction-following ability of LLMs in Hindi. Structured similarly to its English counterpart, it features "verifiable instructions" with heuristically validated responses. These prompts are natively curated by Hindi-proficient specialists to capture local linguistic nuances and Indian cultural context.
  • GSM8K-Hi [^4] https://huggingface.co/datasets/nvidia/GSM8K-Hi: Hindi-GSM8K is the GCP-translated version of the English GSM8K test set. Its samples are meticulously reviewed and corrected by human annotators for quality improvement. These problems typically require 2 to 8 steps to solve, primarily involving elementary arithmetic calculations.
  • MT-Bench-Hi [^5] https://huggingface.co/datasets/nvidia/MT-Bench-Hi: The Hindi MTBench dataset consists of 200 multi-turn prompts designed to evaluate the conversational ability of Hindi LLMs. Eighty percent of its samples are natively created by Hindi specialists, with the remaining 20% translated from the English version, ensuring a balanced and comprehensive evaluation. For this work, we use a subset of 40 samples from MTBench-Hi, focusing on classes such as coding, STEM, math, reasoning, and multiturn interactions. The evaluation is conducted by GPT-4o, with responses rated on a scale of 1-10. It is noted that the scores obtained are on the lower side, as this subset represents areas where Hindi models typically do not excel.
@@ Appendix A Appendix
Given the following sentences: